OpenStudy (anonymous):

Solve this word problem. During the summer you want to earn at least $150 per week. You earn $10 per hour working for a farmer, and you can earn $5 per hour babysitting for your neighbor. You can work at most 25 hours a week. A: Write and solve a system of linear inequalities that models the situation. Let x be the number of hours per week working on the farm and let y be the number of hours per week babysitting. After that it needs to be graphed.

OpenStudy (anonymous):

@zepdrix

OpenStudy (alivejeremy):

Yahoo......... https://answers.yahoo.com/question/index?qid=20130314093742AAy9Hd0

OpenStudy (alivejeremy):

;)

OpenStudy (alivejeremy):

http://www.jiskha.com/display.cgi?id=1236835763

OpenStudy (alivejeremy):

Answer ^^

zepdrix (zepdrix):

\(\large\rm x\) is hours farmed, \(\large\rm y\) is hours babysitting. Total hours must be `less than or equal to 25`.

OpenStudy (anonymous):

Is this it? \[x + y \le 25\]

zepdrix (zepdrix):

good good good c: sorry a little distracted at the moment

OpenStudy (anonymous):

Is that I i have to do ? use that to find out the max number of hours I could work? it sounded like I need more than that.

OpenStudy (anonymous):

I have to earn $150 a week but I can't work more than 25 hours

zepdrix (zepdrix):

We need to create another equation.

zepdrix (zepdrix):

$10 per farm hour, $5 per babysitting hour.

zepdrix (zepdrix):

So the total amount of money made farming is \(\large\rm 10x\). That's the rate $10 per hour, times the number of hours farming x.

zepdrix (zepdrix):

So likewise, the total amount of money made babysitting is \(\large\rm 5y\)

zepdrix (zepdrix):

We need these totals to add up to `at least` 150.

OpenStudy (anonymous):

\[x + y \le 25\] \[10x + 5y \ge 150\]

OpenStudy (anonymous):

are those right?

zepdrix (zepdrix):

Ok great \c:/ So there is our system of equations.

zepdrix (zepdrix):

Then we need to graph this information, ya?

OpenStudy (anonymous):

yep now we just have to find the largest number possible that fit both right?

zepdrix (zepdrix):



OpenStudy (anonymous):

or we go straight to the graph???

zepdrix (zepdrix):

Ya let's just graph :3

OpenStudy (anonymous):

ok. you know what you're doing. I don't. Just going to trust you here.

zepdrix (zepdrix):

\(\large\rm x+y\le25\) Let's get a couple points graphed. First think about it like this: \(\large\rm x+y=25\) If we plug in x=0, \(\large\rm 0+y=25\) Then y=25 solves the equation. This gives us a point, (0,25).

zepdrix (zepdrix):